### v1

@@ 2.2 Orthogonal polynomials
The following theorem can be found in [12, Theorem 2.9].
@@ 3 Orthogonality of GFP of Fibonacci and Lucas polynomials
Horadam, in his paper “New Aspects of Morgan-Voyce polynomials” [9], showed that the Morgan-Voyce polynomials of both types are also Orthogonal polynomials. He used the Chebyshev orthogonal proof to provide a proof and weight function for Morgan-Voyce. In this section, we likewise use the Chebyshev orthogonality proof to provide both a proof for some cases of GFPs and the weight function.
In this section, we present families of GFPs that are orthogonal. Their orthogonality depends on the function $d(x)$ being linear, $g(x)$ being constant, and the choice of a specific weight. However, the question remains as to whether there exist additional families of orthogonal GFPs. This question will be addressed in the following section.
In Section 5, we explore the connections between orthogonal GFPs and probability theory, specifically in the context of random walks.
@@
From [2, Theorem 1.14], we have that when $t=1$ in the previous theorem, the weight $w(x)$ must be unique.
@@ 5.6 Remark
    Indeed, let $\{X_{n},n=0,1,\dots\}$ be a discrete-time birth-and-death chain with $q_{0}=0$ (i.e., 0 is a reflecting state). It follows from [2, Theorem 2.24] that the following statements are equivalent.
@@
    Let $\{X_{n},n=0,1,\dots\}$ be a discrete-time birth-and-death chain with $q_{0}>0$. Then by [2, Theorem 2.32] the following are equivalent.
@@
    Similarly, a corresponding statement applies to continuous-time birth-and-death chains. For the case $\mu_{0}=0$ refer to [2, Theorem 3.42], and for the case $\mu_{0}>0$, see [2, Theorem 3.51].

### v2

@@ 2.2 Orthogonal polynomials
The following theorem can be found in [14, Theorem 2.9].
@@ 3 Orthogonality of GFP of Fibonacci and Lucas polynomials
Horadam, in his paper “New Aspects of Morgan-Voyce Polynomials” [11], showed that the Morgan-Voyce polynomials of both types are orthogonal polynomials. He employed the Chebyshev orthogonality method to derive both a proof and a corresponding weight function for the Morgan-Voyce polynomials. Swamy [16] also proved that $B_{n}$ is orthogonal. André-Jeannin [1, 2] proved that $U_{n}\!\left(\frac{x+p}{2\sqrt{q}}\right)$ and $2q^{n/2}T_{n}\!\left(\frac{x+p}{2\sqrt{q}}\right)$ are orthogonal for real numbers $p$ and $q>0$.
In this section, we likewise apply the Chebyshev orthogonality approach to establish orthogonality for certain cases of GFPs and to determine their associated weight functions.
We present families of GFPs that are orthogonal. Their orthogonality depends on the function $d(x)$ being linear, $g(x)$ being constant, and on the selection of a suitable weight function. However, the question remains as to whether additional families of orthogonal GFPs exist. This issue will be addressed in the following section.
In Section 5, we examine the connections between orthogonal GFPs and probability theory, particularly in the context of random walks.
@@
From [4, Theorem 1.14], we have that when $t=1$ in the previous theorem, the weight $w(x)$ must be unique.
@@ 4 Roots of Generalized Fibonacci Polynomials and their consequences
The roots of some of the polynomials described in Table 1 have already been determined. For example, the roots of the Fibonacci and Lucas polynomials are given in [10, 20], while the roots of the Morgan-Voyce polynomials of the second kind are presented in [16].
André-Jeannin [1, 2] introduced a two-parameter generalization of the Fibonacci polynomials by defining
$U_{n}(p,q;x)=(x+p)\,U_{n-1}(p,q;x)-q\,U_{n-2}(p,q;x),\qquad n\geq 2,$
with $U_{0}(p,q;x)=0$ and $U_{1}(p,q;x)=1$, and
$V_{n}(p,q;x)=(x+p)\,V_{n-1}(p,q;x)-q\,V_{n-2}(p,q;x),\qquad n\geq 2,$
with $V_{0}(p,q;x)=2$ and $V_{1}(p,q;x)=x+p$. André-Jeannin also determined the roots of these polynomials.
@@ 5.6 Remark
    Indeed, let $\{X_{n},n=0,1,\dots\}$ be a discrete-time birth-and-death chain with $q_{0}=0$ (i.e., 0 is a reflecting state). It follows from [4, Theorem 2.24] that the following statements are equivalent.
@@
    Let $\{X_{n},n=0,1,\dots\}$ be a discrete-time birth-and-death chain with $q_{0}>0$. Then by [4, Theorem 2.32] the following are equivalent.
@@
    Similarly, a corresponding statement applies to continuous-time birth-and-death chains. For the case $\mu_{0}=0$ refer to [4, Theorem 3.42], and for the case $\mu_{0}>0$, see [4, Theorem 3.51].